Fermi National Laboratory
 

Accelerator Update for December 4, 2006 - December 6, 2006

 

Monday December 4

The day shift began with Operations monitoring store 5105 (60.81E30) and a Recycler stash (172.3E10), with the Antiproton Source (Pbar) stacking (50.0E10), with Linac, Booster, the Recycler, and the Main Injector (MI) operating normally, with NuMI and MiniBooNE taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

Operations rebooted the MI vacuum VME crate at 9:32 AM. It had stopped responding.

A MI expert began a coalescing study at 10:34 AM.

At 12:02 PM, controls experts began working on a Booster repeater. Pulsed devices and RF were turned off for this work. They had completed their work by 12:16 PM. Booster resumed sending beam.

At 1:29 PM, Operations reported having problems with a Booster BPM program. Operations contacted an expert, but she had already begun work on the problem and had it repaired an hour and a half later.

The MI coalescing study ended and a slipstacking study began at 2:29 PM.

A NuMI expert reported at 2:59 PM that they were back to their normal sump configuration.

At 3:25 PM, Operations reported that beam was falling out of the Main Injector on NuMI only events. The Recycler was stashing during this problem. After the stash ended, NuMI beam made it through MI. Experts investigated.

The Recycler completed its third and last stash of antiprotons for the day shift at 3:30 PM.

The MI slipstacking study ended at 3:49 PM and another coalescing study began.

The MI coalescing study ended at 4 PM.

At 4:17 PM, the NuMI horn tripped off (due to striplines) for the first of two times in forty minutes. It reset both times.

MI experts began a barrier bucket study at 4:34 PM. It ended at 8 PM.

The Recycler completed a stash at 8:49 PM.

Operations terminated store 5105 at 9:33 PM.

Operations put the TeV into shot setup at 9:51 PM.

At 10:54 PM, Operations reported a vacuum burst in the MI-8 line. Many vacuum pumps had tripped off. They all reset okay, but the vacuum around two of the pumps (IP831 and 832) crashed and near-by pumps tripped off again. Operators reset all but those two pumps (those pumps were turned off) and the vacuum recovered.

At 11:14 PM, MI beam began to fall out after injecting the last batch of antiprotons into the TeV for the next store. Operators reduced the intensity to investigate, but the problem disappeared.

Operations established store 5106 at 11:46 PM with an initial luminosity of 208.3E30.

Tuesday December 2

The midnight shift began with Operations monitoring store 5106 (184.59E30) and a stash (15.9E10), with Pbar stacking (57.9E10), and with NuMI and MiniBooNE taking beam.

At 1:52 AM, MIRF8 tripped off. It reset okay.

At 5:39 AM, MIRF12 tripped off. It reset okay.

Operations reported at 8:07 AM that the Recycler had completed three stashes for the midnight shift.

At 10:10 AM, Operations turned off the Proton Source beam so technicians could test the Linac Neutron Therapy Facility (NTF) safety system. During this downtime, experts accessed the Preaccelerator for maintenance, and MI experts worked on their high level RF system.

Techs completed the NTF safety system test at 10:51 AM.

MI RF experts completed their work at 11:10 AM.

The Preaccelerator access ended at 11:45 AM.

Linac resumed sending beam to Pbar, NuMI, and MiniBooNE at 11:54 AM.

At 12:04 PM, Pbar reported having problems with beam positions in the AP1 line.

At 12:12 PM, Operations reported that Booster was having beam problems. Experts were investigating.

Booster resumed sending beam to Pbar, NuMI, and MiniBooNE at 12:29 PM.

A MI expert began a multibatch slipstacking study at 12:57 PM. This study ended at 3:11 PM.

The Recycler completed its second and last stash for the day shift at 3:38 PM.

At 3:39 PM, Operations reported that NuMI beam was again falling out of MI during the stash. After the stash ended, normal beam running resumed.

At 3:42 PM, the NuMI horn tripped off on a relay chassis error. It reset okay.

At 5:07 PM, the Linac Debuncher water skid tripped off. Operators couldn’t reset the skid. Operations called in experts.

A Linac expert managed to turn on the Debuncher water skid at 5:39 PM. He reported that the skid was running warm. Later, operators manually opened the mixing valve to 100% and then after the temperature dropped to its nominal range (96˚F) closed the valve to 75% and then to 55%. The Linac expert bypassed the temperature control loop. He said he would need about four hours to fix the control circuitry.

Linac resumed sending beam at 6:26 PM.

The Recycler completed a stash at 7:40 PM. Operations reported an instrumentation problem (DCCT) during the stash.

At 8:28 PM, a Linac power supply’s (L:QPS215) output began to wander. Operators swapped its control cards with QPS217, but with no luck. They switched back to the original cards. Operators adjusted the supply and resumed normal running, but only for 25 minutes. Operators swapped other control cards.

Operations terminated store 5106 at 10:02 PM.

Operations had D0 ready for an access at 10:37 PM.

At 10:25 PM, Operations reported a problem with a TeV collimator (F491V). The collimator would move to a spot and then stop. Its control card wasn’t responding. Operators tried both a soft and hard reboot, but neither worked. Operations contacted experts who suggested resetting the current position readback. This worked.

At 10:37 PM, the TeV ramp dumped. The TeV sector F3 VCB opened on a transformer temperature indication. An expert investigated and then bypassed the sensor.

Wednesday December 3

The midnight shift began with Operations monitoring the stash (274.0E10), with Pbar stacking (16.6E10), with D0 in access, and with NuMI and MiniBooNE taking beam.

The Recycler completed a stash at 12:57 AM.

At 1:06 AM, D0 reported having trouble with their chiller. An expert came in and fixed the problem (an interlock problem). Unfortunately, D0 said they would need two to three hours to recover their silicon detector.

Operations put the TeV through a dry squeeze at 1:40 AM.

Operations put the TeV into shot setup.

At 2:20 AM, the P1 transfer line beam permit tripped off due to a high current reading from a quadrupole power supply (Q703).

D0 reported that they turned on their silicon detector at 2:43 AM.

Operations established store 5108 at 4:56 AM with an initial luminosity of 178.2E30).

Recycler experts began a study period around 5 AM.

The Plans for Wednesday

The plans for today are to stack, stash, and store. Later today, operators will establish beam to the Switchyard absorber.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 12/06/2006    email Fermilab